10) A home buyer buys a house for $2,155,000. She pays 20% cash, and takes a fixed-rate mortgage for ten years at 7.70% APR. If she makes semi-monthly payments, which of the following is closest to each of her payment?

A) $11,342.47

B) $10,311.34

C) $12,373.61

D) $8249.07
Answer #1

Loan amount = 2,155,000 (1 - 0.2) = 1,724,000

There are 2 payments in a month and there are 12 months in a year. There there are a total of 2 * 12 = 24 periods in a year

Rate = 7.70% / 24 = 0.320833

Number of periods = 10 * 24 = 240

Present value = Payment * [1 - 1 /(1 + r)n]/ r

1,724,000 = Payment * [1 - 1 /(1 + 0.00320833)240]/ 0.00320833

1,724,000 = Payment * [1 - 0.463584] / 0.00320833

1,724,000 = Payment * 167.194619

Payment = $10,311.34